OpenStudy (anonymous):

rectangle ABCD and EFGH are similar, and the ratio of AB to EF is 1:3. in each rectangle, the length is twice the width. the area of ABCD is 32 inches squared. find the length, width and area of EFGH

OpenStudy (anonymous):

width=12 length=24 area=288 #asian swag

OpenStudy (anonymous):

Agreed^

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

the length of EFGH is 24 the width is 12 and the area is 288

OpenStudy (anonymous):

XD :P

OpenStudy (anonymous):

How do u do it though

OpenStudy (anonymous):

A=l*w we know that the length is twice the width: so if x is width then 2x is length. so: A=2x(x) or A=2x^2 A= 32 so 32=2x^2 divide both sides by two 16=x^2 x=4 the width of ABCD is 4 The width of EFGH is three times that of ABCD So the width of EFGH is 12 and the length is twice the width, so the length is 24 12*14=288 288 is the area.

OpenStudy (anonymous):

get it? :)